OpenStudy (anonymous):

I need help figuring out the wording to this Calc 2 problem

OpenStudy (anonymous):

Consider the parametric equation \[x = 16(\cos \theta + \theta \sin \theta)\] \[y = 16( \sin \theta - \theta \cos \theta)\] What is the length of the curve for \[0 \le \theta \le \frac{ 11 }{ 8 }\pi\]

OpenStudy (anonymous):

I understand what the arc length means, but the equations have two separate arcs over the interval

OpenStudy (anonymous):

@mathstudent55 you seem like the most qualified person answering right now, I just need to understand what it's asking for, I can do the math myself

OpenStudy (amistre64):

\[\int_{a}^{b} ds\] given that ds is \[ds=\sqrt{(x')^2+(y')^2}dt\]

OpenStudy (anonymous):

Where does this equation come from? Just so I have a base of understanding. I don't remember seeing anything like that in lecture

OpenStudy (amistre64):

its just the result of takeing the segment of an arc ... and comparing it to a right triangle

OpenStudy (amistre64):



OpenStudy (amistre64):

as we work our way to infinitesimals we get:  and the pythag takes care of the rest